How would you solve this problem, by factoring?
x^3+3x^2-7x=21
thank you!

Answer by jankiz(13)

Set this x^3+3x^2-7x=21 equal to zero by subtractin 21 from both sides. So the equation is x^3+3x^2-7x-21=0. Then find factors that will give you this answer when you use FOIL. Well..... The factors are (x^2-7)(x+3)! Finally, set this set equal to zero to solve for x. So, (x^2-7)(x+3)=0.
Then solve individual set for x. (x^2-7)=0. x^2=7, sqrtx^2=sqrt7, x=sqrt 7 or 2.65. Second set:(x+3)=0, x=-3! You have two solutions for x, but you can only use positive solution when dealing with measurements, etc. Also, you could have used synthetic division to get the same answer, but synthetic division would only have given you the positive solution!